A company is offering a job with a salary of $30,000 for the first year and a 5% raise each year after that. If that 5% raise continues every year, find the amount of money you would earn in a 40-year career.
Answer by Theo(13342)

your first year is 30,000
you earn 5% raise every year.
your growth factor is 1.05
this sounds a lot like a geometric ratio.
An = A1 * r^(n-1)
Sn = A1 * (1 - r^n) / (1-r)
n = 40
A1 = 30,000
A40 = $30,000 * 1.05^39 = $201,142.53
S40 = A1 * (1 - 1.05^40) / (1-1.05) which becomes:
S40 = 30,000 * -6.039988712 / -.05 which becomes:
S40 = -181,199.6614 / -.05 which becomes:
S40 = $3,623,993.227
the individual yearly calculations are shown below:

$$$